Please help me solve this word problem: A pharmacist needs to strengthen a 15% alcohol solution so that it contains 32% alcohol. How much pure alcohol should be added to 100ml of the 15% solution?

Answer by stanbon(75887)

A pharmacist needs to strengthen a 15% alcohol solution so that it contains 32% alcohol. How much pure alcohol should be added to 100ml of the 15% solution?
---------------
15% solution DATA:
Amount = x ml ; Amt of active ingredient = 0.15x ml
-------------------------
Pure alcohol DATA:
Amt = 100-x ml : Amt of active is 100-x ml
--------------------------
32% solution DATA:
Amt = 100 ml ; Amt of active is 0.32*100 = 32 ml
------------------------
EQUATION:
active + active = active
0.15x + 100-x = 32
-0.85x = -68
x = 80 ml
-------------
Amt of pure alcohol needed for the solution = 100-x = 20 ml
=============
Cheers,
Stan H.